OpenStudy (18jonea):

Write the related series for the finite sequence 7.6, 6.3, 5, ... , –1.5. Then evaluate the series. 7.6 + 6.3 + 5 + ... + (–1.5) = 17.4 7.6 + 6.3 + 5 + (–1.5) = 17.4 7.6 + 6.3 + 5 + ... + (–1.5) = 24.4 none of these

OpenStudy (18jonea):

@Michele_Laino

OpenStudy (18jonea):

@tkhunny

OpenStudy (michele_laino):

it is an arithmetic series, whose constant is: \[\huge d = 7.6 - 6.3 = 6.3 - 5 = ...?\]

OpenStudy (18jonea):

what does that mean?

OpenStudy (michele_laino):

what is \(d\) ?

OpenStudy (18jonea):

so would you find the value of 7.6- 6.3 amd 6.3-5

OpenStudy (18jonea):

it is 1.3

OpenStudy (michele_laino):

yes! More precisely it is \(d=-1.3\)

OpenStudy (michele_laino):

now, the last term is: \(-1.5\), so we can write: \[\huge - 1.5 = 7.6 + \left( {n - 1} \right) \cdot \left( { - 1.3} \right)\] where \(n\) is the number of terms of the sequence. Please solve for \(n\)

OpenStudy (18jonea):

how would i solve for N?

OpenStudy (michele_laino):

here is my hint: we can write this: \[\huge n - 1 = \frac{{ - 1.5 - 7.6}}{{ - 1.3}} = ...?\] please continue

OpenStudy (18jonea):

-7

OpenStudy (michele_laino):

I got \(n-1=7\)

OpenStudy (18jonea):

o yeah sorry two negatives make a positive

OpenStudy (18jonea):

so n=8

OpenStudy (michele_laino):

that's right! the number of terms of the sequence is: \[\huge n =8\]

OpenStudy (michele_laino):

now, the sum of all terms of the sequence is given by the subsequent formula: \[\large {S_8} = \frac{{{a_1} + {a_8}}}{2} \times 8 = \frac{{7.6 + \left( { - 1.5} \right)}}{2} \times 8 = ...?\]

OpenStudy (18jonea):

i got 24.4

OpenStudy (michele_laino):

that's right!

OpenStudy (18jonea):

so c

OpenStudy (michele_laino):

yes!

OpenStudy (18jonea):

Thank you so much for your help

OpenStudy (michele_laino):

:)